OpenStudy (anonymous):

The height of a triangle is 5 centimeters less than the base. The area of the triangle 3 square centimeters. Find the length of the base and height of the triangle.

OpenStudy (anonymous):

Do you know the formula of the area of a triangle?

OpenStudy (anonymous):

1/2 base times height right?

OpenStudy (anonymous):

Yup. So you know that height of the triangle is 5 cm less than the base. A way to represent this in equation form would be as so:\[height = base - 5cm\] you know that your equation of a triangle is:\[Area = \frac{ 1 }{ 2 }(base)(height)\]you plug what you have as your height into the formula and also the known area (which is 3 cm^2):\[3 cm^2= (base)(base-5cm)\]can you solve for the base from here?

OpenStudy (anonymous):

oops I meant this as the final equation for the solution of the base:\[3cm^2 = \frac{ 1 }{ 2 }(base)(base-5cm)\]ignore the top one

OpenStudy (anonymous):

Thank you!